### v3

@@ Abstract
Many small to large organizations have adopted the Microservices Architecture (MSA) style to develop and deliver their core businesses. Despite the popularity of MSA in the software industry, there is a limited evidence-based and thorough understanding of the types of issues (e.g., errors, faults, failures, and bugs) that microservices system developers experience, the causes of the issues, and the solutions as potential fixing strategies to address the issues. To ameliorate this gap, we conducted a mixed-methods empirical study that collected data from 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, 15 interviews, and an online survey completed by 150 practitioners from 42 countries across 6 continents. Our analysis led to comprehensive taxonomies for the issues, causes, and solutions. The findings of this study inform that Technical Debt, Continuous Integration and Delivery, Exception Handling, Service Execution and Communication, and Security are the most dominant issues in microservices systems. Furthermore, General Programming Errors, Missing Features and Artifacts, and Invalid Configuration and Communication are the main causes behind the issues. Finally, we found 177 types of solutions that can be applied to fix the identified issues. Based on our study results, we formulated future research directions that could help researchers and practitioners to engineer emergent and next-generation microservices systems.
@@ 1 Introduction
The software industry has recently witnessed the growing popularity of the Microservices Architecture (MSA) style as a promising design approach to develop applications that consist of multiple small, manageable, and independently deployable services [1]. Software development organizations may have adopted or planned to use the MSA style for various reasons. Specifically, some of them want to increase the scalability of applications using the MSA style, while others use it to quickly release new products and services to the customers, whereas it is argued that the MSA style can also help build autonomous development teams [2, 3]. From an architectural perspective, a microservices system (a system that adopts the MSA style) entails a significant degree of complexity both at the design phase as well as at runtime configuration [4]. This implies that the MSA style brings unique challenges for software organizations, and many quality attributes may be (positively or negatively) influenced [5]. For example, service level security may be impacted because microservices are developed and deployed by various technologies (e.g., Docker containers [6]) and tools that are potentially vulnerable to security attacks [4, 7]. Data management is also influenced because each microservice needs to own its domain data and logic [8]. This can, for example, challenge achieving and managing data consistency across multiple microservices.
@@
Zimmermann argues that MSA is not entirely new from Service-Oriented Architecture (SOA) (e.g., “microservices constitute one particular implementation approach to SOA – service development and deployment”) [9]. Similarly, Márquez and Astudillo discovered that some existing design rationale and patterns from SOA fit the context for MSA [10]. However, an important body of literature (e.g., [11, 3]) has concluded that there are overwhelming differences between microservices systems, monolithic systems, and traditional service-oriented systems in terms of design, implementation, testing, and deployment. Gupta and Palvankar indicated that even having SOA experience can lead to suboptimal decisions (e.g., excessive service calls) in microservices systems [12]. Hence, microservices systems may have an additional and specific set of issues. Borrowing the idea from [13], we define issues in this study as errors, faults, failures, and bugs that occur in a microservices system and consequently impact its quality and functionality. Hence, there is a need to leverage existing methods or derive new practices, techniques, and tools to address the specific and additional issues in microservices systems.
@@
Recently, a number of studies have investigated particular issues (e.g., code smell [14], debugging [15], performance [16]) in microservices systems. Despite these efforts, there is no in-depth and comprehensive study on the nature of different types of issues that microservices developers face, the potential causes of these issues, and possible fixing strategies for these issues. Jamshidi et al. believed that this can be partially attributed to the fact that researchers have limited access to industry-scale microservices systems [3]. The empirical knowledge on the nature of issues occurring in microservices systems can be useful from the following perspectives: (i) understanding common issues in the design and implementation of microservices systems and how to avoid them, (ii) identifying trends in the types of issues that arise in microservices systems and how to address them effectively, (iii) experienced microservices developers can be allocated to address the most frequent and challenging issues, (iv) novice microservices developers can quickly be informed of empirically-justified issues and avoid common mistakes, and (v) the industry and academic communities can synergies theory and practice to develop tools and techniques for the frequently reported issues in microservices systems.
@@
This work aims to systematically and comprehensively study and categorize the issues that developers face in developing microservices systems, the causes of the issues, and the solutions (if any). To this end, we conducted a mixed-methods empirical study following the guideline proposed by Easterbrook and his colleagues [17]. We collected the data by (i) mining 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, (ii) conducting 15 interviews, and (iii) deploying an online survey completed by 150 practitioners to develop the taxonomies of the issues, their causes and solutions in microservices systems. The key findings of this study are:
@@
This paper has extended our previous work [18] by adding two new research questions (RQ3 and RQ4) and expanding and enhancing the results of RQ1 and RQ2 with increased volume and variety of data and a mixed-methods approach. Specifically, we explored 10 more open-source microservices systems on GitHub (now 15 projects), interviewed 15 practitioners, and conducted an online survey with 150 microservices practitioners for getting their perspectives on the proposed taxonomies of the issues, their causes and solutions in microservices systems, as well as the mapping between the issues, causes, and solutions.
@@
    We made the dataset of this study available online [19], which includes the data collection and analysis from GitHub and microservices practitioners, as well as detailed hierarchies of the taxonomies of issues, causes, and solutions, to enable replication of this study and conduct future research.
The remainder of the paper is structured as follows. Section 2 details the research methodology employed. Section 3 presents the results of our study. Section 4 discusses the relationship between the issues, causes, and solutions, along with the implications and the threats to the validity of our results. Section 6 reviews related work, and Section 7 draws conclusions and outlines avenues for future work.
@@ 2.2.1 Step A – Identify & Select MSA-based OSS Projects
The specified RQs require us to identify and select MSA-based OSS projects, representing a repository of developer discussions and knowledge, to extract the issues, their causes, and solutions. This means that the RQs guided the development of search strings based on the recommendations and steps from [20] for string composition to retrieve developer discussions on microservices projects deployed on GitHub [21]. We formulated the search string using the format [keyword-1 [OR logic] keyword-2 … [OR logic] … keyword-N], where keywords represented the synonyms as [‘micro service’ OR ‘micro-service’ OR ‘microservice’ OR ‘Micro service’ OR ‘Micro-service’ OR ‘Microservice’]. To extract MSA issues, we selected GitHub, which is one of the most popular and rapidly growing platforms for social coding and community-driven collaborative development of OSS. GitHub represents a modern genre of software forges that unifies traditional methods of development (e.g., version control, code hosting) with features of socio-collaborative development (e.g., issue tracking, pull requests) [22]. The variety and magnitude of OSS available on GitHub also inspired our choice to investigate the largest OSS platform in the world, with approximately 40 million users and 28 million publicly available project repositories.
Based on the search string, we searched for the title and description of the OSS projects deployed on the GHTorrent dump hosted on Google Cloud. The search helped us retrieve a total of 2,690 potentially relevant MSA-based OSS projects for investigation. To shortlist and eventually select the projects pertinent to the outlined RQs, we applied multi-criteria filtering [23], considering a multitude of aspects such as the popularity or perceived significance of a project in the developers’ community (represented as total stars), adoption by or interests of developers (total forks), and the total number of developers involved (total contributors) for the project. As shown in Figure 2 (Step B), we only selected the projects that have (i) more than 10 stars and forks, (ii) an English description, and (iii) three or more contributors. This led to a shortlist of 167 microservices projects. To eliminate the instances of potential false positives, i.e., avoiding bias in construct validity, such as misleading project names and mockup code, we contacted the top three contributors of each project via their emails to clarify about:
@@ 2.2.2 Step B – Synthesize Issues, Causes, and Solutions
Raw Data Collection: We chose 15 microservices projects (see Table 1) as the source for building the dataset to answer our RQs. These 15 projects were chosen because they are significantly larger than other microservices projects hosted on GitHub. Hence, it is highly likely that their contributors had more discussions about the type of issues, causes, and solutions in issue tracking systems. The discussions relating to a software system can usually be captured in issue tracking systems [24]. We initially extracted 10,222 issue titles, issue links, issue opening and closing dates, and the number of contributors for each issue through our customized Python script (see the Raw Data sheet in [19]). We stored this information in MySQL and exported it into MS Excel sheets for further processing. We extracted only closed issues because it could increase the chances of answering all our RQs (e.g., solutions).
Issues Screening: The first author further scanned the 10,222 issues to check if an issue has been closed or is still open. All open issues were discarded because an open issue is ongoing with many of its causes unknown and most likely solution(s) not found. Furthermore, after selecting only the closed issues, the first author further eliminated (i) issues without a detailed description, (ii) general questions, opinions, feedback, and ideas, (iii) feature requests (e.g., enhancements, proposals), (iv) announcements (e.g., about new updates), (v) duplicated issues, (vi) issues that had only one participant, and (vii) stale issues. After this step, we got 5,115 issues (see the Selected Issues (Round 1) sheet in [19]). We had second round of screening on these 5,115 issues to check whether these issues are related to our RQs or not, and after comprehensively analyzing them we found 2,641 issues that were related to our RQs (see the Selected Issues (Round 2) sheet in [19]).
@@
Data Extraction: Issues, causes, solutions, and their corresponding data were extracted based on the guidelines for mining software engineering data from GitHub [25]. We extracted issue titles, issue links, the number of participants for each issue, key points of issues, key points of causes, and key points of solutions (see the Raw Data and Initial Codes sheet in [19]). Data such as issue title, issue link, and number of participants for each issue were used to capture general information about developer discussions. On the other hand, key points of issues, causes, and solutions were used to document the data for answering RQ1, RQ2, RQ3.
Data Analysis: To synthesize the issues, we used the thematic analysis approach [26] to identify the categories of issues, causes, and solutions. The thematic analysis approach is composed of five steps. (i) Familiarizing with data: The first author repeatedly read the project’s contributor’s discussion and documented all discussed key points about issues, causes, and solutions. (ii) Generating initial codes: after data familiarization, the first author generated an initial list of codes from the extracted data (see the Initial Codes sheet in [19]). (iii) Searching for the types of issues, causes, and solutions: The first and second authors analyzed the initially generated codes and brought them under the specific types of issues,causes, and solutions. (iv) Reviewing types of issues, causes, and solutions: All the authors reviewed and refined the coding results with the corresponding types of issues, causes, and solutions. We separated, merged, and dropped several issues, causes, and solutions based on a mutual discussion between all the authors. (v) Defining and naming categories: We defined and further refined all the types of issues, causes, and solutions under precise and clear subcategories and categories (see Figure 4). Throughout this process, we engaged in the collaborative refinement and organization of issues, causes, and solutions into specific subcategories and categories. The naming and assignment of issues, causes, and solutions in microservices systems to the appropriate subcategories were conducted through consensus among all the authors, ensuring the development of a comprehensive and coherent taxonomy. Finally, we introduced three levels of categories for managing the identified issues, causes, and solutions. First, we organized the types of issues, causes, and solutions under a specific subcategory (e.g., service dependency in Service Design Debt). Then we arranged the subcategories under a specific category (e.g., Service Design Debt in Technical Debt).
@@ 2.3.1 Preparing a Protocol
The first author conducted 15 online interviews with microservices practitioners through Zoom, Tencent Meeting, and Microsoft Teams. Before conducting actual interviews, we also conducted two pilot interviews with microservices practitioners to check the understandability and comprehensiveness of interview questions. However, we did not include their answers in our dataset. In total, we conducted 15 actual interviews, and each interview took 35-45 minutes. It is argued that conducting 12 to 15 interviews with homogeneous groups is enough to reach saturation [27]. In our case, after completing 15 interviews, we noticed a clear saturation in the responses to our interview questions, indicating that we had gathered enough information to address our interview objectives effectively. Therefore, we decided not to conduct any additional interviews. We conducted semi-structured interviews based on an interview guide (see the Interview Questionnaire sheet in [19]), which contains closed-ended questions related to participant demographics and open-ended questions related to general categories of microservices issues, causes, and solutions.
@@ 2.3.3 Data Analysis
We applied a thematic analysis method [28] to analyze the recorded interviews. Before applying the thematic analysis method, the first author prepared the text transcripts from audio recordings. The first author read the interview transcripts and coded them using the MAXQDA tool. We dropped several sentences unrelated to “microservices issues, causes, and solutions”. After removing the extraneous information from the transcribed interviews, the first author coded the interview transcripts’ contents to get the answers to the interview questions. To ensure the quality of the codes, the second author verified the initial codes created by the first author and provided suggestions for improvement. After the initial coding and subsequent enhancements, all the authors collaborated in analyzing the codes. Conflicts in the codes were addressed through collaborative discussions and brainstorming sessions. Finally, we generated a total of 28 types of issues (classified into 15 subcategories and 11 categories), 28 types of causes (classified into 15 subcategories and 7 categories), and 30 types of solutions (classified into 4 subcategories and 3 categories). During the interviews, we got 48 instances of issues, 31 instances of causes, and 36 instances of solutions. Among these instances, we identified 14 types of issues, 21 types of causes, and 23 types of solutions that were not part of the taxonomies we derived from the 15 open-source microservices systems. Later, we exported the analyzed interview data from the MAXQDA tool to an MS Excel sheet (i.e., the Interview Results sheet in [19]) to make the part of taxonomies of issues, causes, and solutions in microservices systems from the interview data.

### v4

@@ Abstract
Many small to large organizations have adopted the Microservices Architecture (MSA) style to develop and deliver their core businesses. Despite the popularity of MSA in the software industry, there is a limited evidence-based and thorough understanding of the types of issues (e.g., errors, faults, failures, and bugs) that microservices system developers experience, the causes of the issues, and the solutions as potential fixing strategies to address the issues. To ameliorate this gap, we conducted a mixed-methods empirical study that collected data from 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, 15 interviews, and an online survey completed by 150 practitioners from 42 countries across 6 continents. Our analysis led to comprehensive taxonomies for the issues, causes, and solutions. The findings of this study inform that Technical Debt, Continuous Integration and Delivery, Exception Handling, Service Execution and Communication, and Security are the most dominant issues in microservices systems. Furthermore, General Programming Errors, Missing Features and Artifacts, and Invalid Configuration and Communication are the main causes behind the issues. Finally, we found 177 types of solutions that can be applied to fix the identified issues. Based on our study results, we propose a future research framework that outlines key problem dimensions and actionable study strategies to support the engineering of emergent and next-generation microservices systems.
@@ 1 Introduction
The software industry has recently witnessed the growing popularity of the Microservices Architecture (MSA) style as a promising design approach to develop applications that consist of multiple small, manageable, and independently deployable services [fowler2014microservices]. Software development organizations may have adopted or planned to use the MSA style for various reasons. Specifically, some of them want to increase the scalability of applications using the MSA style, while others use it to quickly release new products and services to the customers, whereas it is argued that the MSA style can also help build autonomous development teams [Davide2017Processes, jamshidi2018microservices]. From an architectural perspective, a microservices system (a system that adopts the MSA style) entails a significant degree of complexity both at the design phase as well as at runtime configuration [newman2020building]. This implies that the MSA style brings unique challenges for software organizations, and many quality attributes may be (positively or negatively) influenced [waseem2020systematic]. For example, service level security may be impacted because microservices are developed and deployed by various technologies (e.g., Docker containers [combe2016docker]) and tools that are potentially vulnerable to security attacks [newman2020building, yu2019survey]. Data management is also influenced because each microservice needs to own its domain data and logic [wagner2018net]. This can, for example, challenge achieving and managing data consistency across multiple microservices.
@@
Zimmermann argues that MSA is not entirely new from Service-Oriented Architecture (SOA) (e.g., “microservices constitute one particular implementation approach to SOA – service development and deployment”) [zimmermann2017microservices]. Similarly, Márquez and Astudillo discovered that some existing design rationale and patterns from SOA fit the context for MSA [1-marquez2018actual]. However, an important body of literature (e.g., [esposito2016challenges, jamshidi2018microservices]) has concluded that there are overwhelming differences between microservices systems, monolithic systems, and traditional service-oriented systems in terms of design, implementation, testing, and deployment. Gupta and Palvankar indicated that even having SOA experience can lead to suboptimal decisions (e.g., excessive service calls) in microservices systems [Dinkar2020Pitfalls]. Hence, microservices systems may have an additional and specific set of issues. Borrowing the idea from [ren2020understanding], we define issues in this study as errors, faults, failures, and bugs that occur in a microservices system and consequently impact its quality and functionality. Hence, there is a need to leverage existing methods or derive new practices, techniques, and tools to address the specific and additional issues in microservices systems.
@@
Recently, a number of studies have investigated particular issues (e.g., code smell [taibi2018definition], debugging [ZhouIEEE], performance [WuNoms]) in microservices systems. Despite these efforts, there is no in-depth and comprehensive study on the nature of different types of issues that microservices developers face, the potential causes of these issues, and possible fixing strategies for these issues. Jamshidi et al. believed that this can be partially attributed to the fact that researchers have limited access to industry-scale microservices systems [jamshidi2018microservices]. The empirical knowledge on the nature of issues occurring in microservices systems can be useful from the following perspectives: (i) understanding common issues in the design and implementation of microservices systems and how to avoid them, (ii) identifying trends in the types of issues that arise in microservices systems and how to address them effectively, (iii) experienced microservices developers can be allocated to address the most frequent and challenging issues, (iv) novice microservices developers can quickly be informed of empirically-justified issues and avoid common mistakes, and (v) the industry and academic communities can synergies theory and practice to develop tools and techniques for the frequently reported issues in microservices systems.
@@
This work aims to systematically and comprehensively study and categorize the issues that developers face in developing microservices systems, the causes of the issues, and the solutions (if any). To this end, we conducted a mixed-methods empirical study following the guideline proposed by Easterbrook and his colleagues [easterbrook2008selecting]. We collected the data by (i) mining 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, (ii) conducting 15 interviews, and (iii) deploying an online survey completed by 150 practitioners to develop the taxonomies of the issues, their causes and solutions in microservices systems. The key findings of this study are:
@@
This paper has extended our previous work [waseem2021nature] by adding two new research questions (RQ3 and RQ4) and expanding and enhancing the results of RQ1 and RQ2 with increased volume and variety of data and a mixed-methods approach. Specifically, we explored 10 more open-source microservices systems on GitHub (now 15 projects), interviewed 15 practitioners, and conducted an online survey with 150 microservices practitioners for getting their perspectives on the proposed taxonomies of the issues, their causes and solutions in microservices systems, as well as the mapping between the issues, causes, and solutions.
@@
    We proposed a future research framework that synthesizes the major problem dimensions and outlines actionable study strategies (Table LABEL:tab:future_framework), thereby translating our empirical findings into a concrete research agenda for the microservices community.
  4. 4.
    We made the dataset of this study available online [replpack], which includes the data collection and analysis from GitHub and microservices practitioners, as well as detailed hierarchies of the taxonomies of issues, causes, and solutions, to enable replication of this study and conduct future research.
The remainder of the paper is structured as follows. Section 2 details the research methodology employed. Section 3 presents the results of our study. Section 4 discusses the relationship between the issues, causes, and solutions, along with the implications and evidence-driven future research agenda. Section 5 clarifies the threats to the validity of our results. Section 6 reviews related work, and Section 7 draws conclusions and outlines avenues for future work.
@@ 2.2.1 Step A – Identify & Select MSA-based OSS Projects
The specified RQs require us to identify and select MSA-based OSS projects, representing a repository of developer discussions and knowledge, to extract the issues, their causes, and solutions. This means that the RQs guided the development of search strings based on the recommendations and steps from [SystematicSearchMap2018] for string composition to retrieve developer discussions on microservices projects deployed on GitHub [surana2020tool]. We formulated the search string using the format [keyword-1 [OR logic] keyword-2 … [OR logic] … keyword-N], where keywords represented the synonyms as [‘micro service’ OR ‘micro-service’ OR ‘microservice’ OR ‘Micro service’ OR ‘Micro-service’ OR ‘Microservice’]. To extract MSA issues, we selected GitHub, which is one of the most popular and rapidly growing platforms for social coding and community-driven collaborative development of OSS. GitHub represents a modern genre of software forges that unifies traditional methods of development (e.g., version control, code hosting) with features of socio-collaborative development (e.g., issue tracking, pull requests) [kalliamvakou2016depth]. The variety and magnitude of OSS available on GitHub also inspired our choice to investigate the largest OSS platform in the world, with approximately 40 million users and 28 million publicly available project repositories.
Based on the search string, we searched for the title and description of the OSS projects deployed on the GHTorrent dump hosted on Google Cloud. The search helped us retrieve a total of 2,690 potentially relevant MSA-based OSS projects for investigation. To shortlist and eventually select the projects pertinent to the outlined RQs, we applied multi-criteria filtering [GitHubFilter2020], considering a multitude of aspects such as the popularity or perceived significance of a project in the developers’ community (represented as total stars), adoption by or interests of developers (total forks), and the total number of developers involved (total contributors) for the project. As shown in Figure 2 (Step B), we only selected the projects that have (i) more than 10 stars and forks, (ii) an English description, and (iii) three or more contributors. This led to a shortlist of 167 microservices projects. To eliminate the instances of potential false positives, i.e., avoiding bias in construct validity, such as misleading project names and mockup code, we contacted the top three contributors of each project via their emails to clarify about:
@@ 2.2.2 Step B – Synthesize Issues, Causes, and Solutions
Raw Data Collection: We chose 15 microservices projects (see Table 1) as the source for building the dataset to answer our RQs. These 15 projects were chosen because they are significantly larger than other microservices projects hosted on GitHub. Hence, it is highly likely that their contributors had more discussions about the type of issues, causes, and solutions in issue tracking systems. The discussions relating to a software system can usually be captured in issue tracking systems [Viviani21]. We initially extracted 10,222 issue titles, issue links, issue opening and closing dates, and the number of contributors for each issue through our customized Python script (see the Raw Data sheet in [replpack]). We stored this information in MySQL and exported it into MS Excel sheets for further processing. We extracted only closed issues because it could increase the chances of answering all our RQs (e.g., solutions).
Issues Screening: The first author further scanned the 10,222 issues to check if an issue has been closed or is still open. All open issues were discarded because an open issue is ongoing with many of its causes unknown and most likely solution(s) not found. Furthermore, after selecting only the closed issues, the first author further eliminated (i) issues without a detailed description, (ii) general questions, opinions, feedback, and ideas, (iii) feature requests (e.g., enhancements, proposals), (iv) announcements (e.g., about new updates), (v) duplicated issues, (vi) issues that had only one participant, and (vii) stale issues. After this step, we got 5,115 issues (see the Selected Issues (Round 1) sheet in [replpack]). We had second round of screening on these 5,115 issues to check whether these issues are related to our RQs or not, and after comprehensively analyzing them we found 2,641 issues that were related to our RQs (see the Selected Issues (Round 2) sheet in [replpack]).
@@
Data Extraction: Issues, causes, solutions, and their corresponding data were extracted based on the guidelines for mining software engineering data from GitHub [gousios2017mining]. We extracted issue titles, issue links, the number of participants for each issue, key points of issues, key points of causes, and key points of solutions (see the Raw Data and Initial Codes sheet in [replpack]). Data such as issue title, issue link, and number of participants for each issue were used to capture general information about developer discussions. On the other hand, key points of issues, causes, and solutions were used to document the data for answering RQ1, RQ2, RQ3.
Data Analysis: To synthesize the issues, we used the thematic analysis approach [cruzes2011recommended] to identify the categories of issues, causes, and solutions. The thematic analysis approach is composed of five steps. (i) Familiarizing with data: The first author repeatedly read the project’s contributor’s discussion and documented all discussed key points about issues, causes, and solutions. (ii) Generating initial codes: after data familiarization, the first author generated an initial list of codes from the extracted data (see the Initial Codes sheet in [replpack]). (iii) Searching for the types of issues, causes, and solutions: The first and second authors analyzed the initially generated codes and brought them under the specific types of issues,causes, and solutions. (iv) Reviewing types of issues, causes, and solutions: All the authors reviewed and refined the coding results with the corresponding types of issues, causes, and solutions. We separated, merged, and dropped several issues, causes, and solutions based on a mutual discussion between all the authors. (v) Defining and naming categories: We defined and further refined all the types of issues, causes, and solutions under precise and clear subcategories and categories (see Figure 4). Throughout this process, we engaged in the collaborative refinement and organization of issues, causes, and solutions into specific subcategories and categories. The naming and assignment of issues, causes, and solutions in microservices systems to the appropriate subcategories were conducted through consensus among all the authors, ensuring the development of a comprehensive and coherent taxonomy. Finally, we introduced three levels of categories for managing the identified issues, causes, and solutions. First, we organized the types of issues, causes, and solutions under a specific subcategory (e.g., service dependency in Service Design Debt). Then we arranged the subcategories under a specific category (e.g., Service Design Debt in Technical Debt).
@@ 2.3.1 Preparing a Protocol
The first author conducted 15 online interviews with microservices practitioners through Zoom, Tencent Meeting, and Microsoft Teams. Before conducting actual interviews, we also conducted two pilot interviews with microservices practitioners to check the understandability and comprehensiveness of interview questions. However, we did not include their answers in our dataset. In total, we conducted 15 actual interviews, and each interview took 35-45 minutes. It is argued that conducting 12 to 15 interviews with homogeneous groups is enough to reach saturation [guest2006many]. In our case, after completing 15 interviews, we noticed a clear saturation in the responses to our interview questions, indicating that we had gathered enough information to address our interview objectives effectively. Therefore, we decided not to conduct any additional interviews. We conducted semi-structured interviews based on an interview guide (see the Interview Questionnaire sheet in [replpack]), which contains closed-ended questions related to participant demographics and open-ended questions related to general categories of microservices issues, causes, and solutions.
@@ 2.3.3 Data Analysis
We applied a thematic analysis method [braun2006using] to analyze the recorded interviews. Before applying the thematic analysis method, the first author prepared the text transcripts from audio recordings. The first author read the interview transcripts and coded them using the MAXQDA tool. We dropped several sentences unrelated to “microservices issues, causes, and solutions”. After removing the extraneous information from the transcribed interviews, the first author coded the interview transcripts’ contents to get the answers to the interview questions. To ensure the quality of the codes, the second author verified the initial codes created by the first author and provided suggestions for improvement. After the initial coding and subsequent enhancements, all the authors collaborated in analyzing the codes. Conflicts in the codes were addressed through collaborative discussions and brainstorming sessions. Finally, we generated a total of 28 types of issues (classified into 15 subcategories and 11 categories), 28 types of causes (classified into 15 subcategories and 7 categories), and 30 types of solutions (classified into 4 subcategories and 3 categories). During the interviews, we got 48 instances of issues, 31 instances of causes, and 36 instances of solutions. Among these instances, we identified 14 types of issues, 21 types of causes, and 23 types of solutions that were not part of the taxonomies we derived from the 15 open-source microservices systems. Later, we exported the analyzed interview data from the MAXQDA tool to an MS Excel sheet (i.e., the Interview Results sheet in [replpack]) to make the part of taxonomies of issues, causes, and solutions in microservices systems from the interview data.